# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

### v1

@@ IV-B2 Data Generation and Completion
Hyperspectral images (HSI) can provide richer spectral spatial information. Meanwhile, the acquisition cost of HSI is much higher than that of multispectral images. Therefore, subsequent scientists then used easily available multispectral remote sensing images to generate hyperspectral images [91, 92]. However, the hyperspectral maps generated by the diffusion model used need to match the input noise dimensions with the spectral bands, which leads to too much noise sampling space, thus making the model convergence difficult. To solve this problem, Zhang et al [91] proposed a spectral folding technique to convert the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al [92] used Conditional Vector Quantization GAN (VQGAN) [93] to obtain the Latent Diffusion Model latententent space for the hyperspectral map, and performed the diffusion model training and sampling in this space process. Afterwards, in order to address the status quo of high cost and low spatial resolution, a DM-based hyperspectral spectral decomposition method, DiffUn [94], was proposed by Deng et al. The spectral prior distributions learned by unconditional DM and the likelihoods derived from the LMM were combined to facilitate the deblurring performance and synthesize the hyperspectral maps.A novel ISPDiff [95] model was proposed by Dong et al. A transparent unified scale-flexible model was designed to solve such problems using an iterative optimization algorithm with MSRN for scale-matching recovery and CMUN for cross-scale upsampling.
@@
Overall, image-to-image generation methods based on conditional diffusion models [67, 68], which use bootstrap images as conditional inputs to conditional DDPM, are essentially generating images from noise rather than directly. In order to realize a true image-to-image conversion, Yu et al. proposed UnmixDiff [96] to address the feasibility of depth generation model in HSI synthesis task, which synthesizes HSI from the perspective of material distribution and introduces a clear physical interpretation into the hyperspectral image synthesis task. By introducing an unmixing autoencoder in the diffusion generation model, the hyperspectral image synthesis is projected from the high-dimensional spectral domain to the low-dimensional abundance domain, which greatly reduces the computational complexity and effectively improves the spectral confidence of the generated hyperspectral images. By constructing a diffusion generation model in the abundance space, at the same time, hsi with real spatial details can be generated to maintain spectral consistency. Also at the same time this type of approach, which first employs self-supervised learning to extract the feature embeddings of the input patches, and then uses these embeddings as conditional inputs to guide the learning of the diffusion model. Notably, this preserves the spatial arrangement of each color patch in the original image so that the generated patches can be assembled into a large and coherent image.
@@ IV-B3 Classification
Among the applications in the hyperspectral field, classification task is very important in image analysis as it can accurately identify and characterize different structures and anomalies in an image. It helps professionals in their ability to interpret large amounts of complex data Despite this potential, image classification aims at obtaining more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98]. However, hyperspectral images contain many diverse and complex scenes and information, which increases the difficulty of accurate classification. Therefore, the current use of diffusion model to enhance classification is still a major challenge that needs to be further strengthened and solved. Meanwhile, because diffusion models can learn and simulate more complex data distributions than other deep learning models, more and more people are committed to applying diffusion models to hyperspectral image classification, which leads to better and better classification results [99, 100, 101, 102, 103, 104].
For image classification, hyperspectral images exhibit more complex data distributions, posing greater challenges to the application of traditional deep learning models in classification. The good thing, however, is that diffusion models can better capture the spectral-spatial joint features of hyperspectral images [105, 106, 107], improve the classification accuracy [108, 109], and improve the generalization ability [110]. Based on the diffusion model to realize the above properties, Zhou et al [111] used the temporal information of the diffusion model to build a time-stepping feature library, and also proposed a dynamic fusion module to combine spectral-spatial features with temporal features, and these improvements can make it possible to obtain sufficient image information before classification.Li et al [112] designed a two-branched diffusion model, which extracts features from HSI and LiDAR images, which achieves complementarity of information between different modal images, enhances the distinguishability between pixels, and also shows excellent performance in more complex hyperspectral tasks [37].Qu et al [113] go deeper by setting the encoder of the two-branch diffusion model to operate in parameter-sharing mode, which ensures that the shared features are extracted from the multimodal images. In addition, Chen et al [114] utilized the diffusion model to assist deep subspace construction and achieved excellent HSI classification performance in an unsupervised manner. Different from the above approaches, Qu et al [115] used a diffusion model to combine the SR (super-resolution) network and classification network to progressively reconstruct high-quality images by iteratively iterating the image super-resolution and classification tasks to achieve mutual guidance of the two tasks. The multi-scale classification results are utilized to guide the target-scale image to learn new knowledge and further reconstruct high-quality images. The super-resolution performance and classification performance are significantly improved.Jiang et al. proposed an image fusion [107] classification network based on Denoising Diffusion Probabilistic Models (DDPM), and after fusion, the classification accuracy is significantly improved.Zhu et al. proposed a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN) based on Denoising Diffusion Probabilistic Models (DDPM) [102], which is structured spatially self-attentive denoising module (SSAD) and spectral group self-attentive denoising module (SGSAD), and improves the efficiency of spectral spatial feature learning by designing new DDPM models loss and CL with log absolute error (LAE), as well as designing an adaptive weighted addition module (AWAM) and cross-time step spectral spatial fusion module (CTSSFM), which can effectively mitigate the interference due to spatial spectral heterogeneity and noise effects, etc., to effectively mitigate the interference due to spatial spectral heterogeneity and better classification.
@@ IV-B4 Segmentation
Image segmentation is an important task in computer vision, like semantic segmentation [116], image segmentation [14, 117], and in the field of deep learning, semantic segmentation is a subtask of image segmentation [112, 118]. Hyperspectral image segmentation, which investigates simplifying the complexity of an image by decomposing it into multiple meaningful image segments. Specifically, it facilitates image analysis by providing useful information about hyperspectral related fields. However, deep learning models typically require a large number of different pixel-annotated training data to produce generalizable results. Nonetheless, the number of images and labels that can be used for hyperspectral image segmentation is limited due to the time, cost, and expertise required. For this reason, diffusion modeling has emerged as a promising approach in image segmentation research that synthesizes labeled data and eliminates the need for pixel annotation data [102, 104].
@@
Dong et al [95] are using Denoising Diffusion Probabilistic Models combined with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks.Chen et al [119] design image conditional mixing guides and combine them with Denoising Diffusion Probabilistic Models.Zhang et al [110] utilize Diffusion Probabilistic Models for labeling information.Zhu et al [120] utilize Variational Auto-Encoder as well as semi-supervised diffusion model for labeling annotation.Li et al. proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to improve the pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism.Zhu et al. proposed a diffusion-based diffusion model coupled with contrast learning for diffusion contrast representation networks (DiffCRNs) [102] to extract the spatial-spectral features and finally generates pixel-level classification circles. Xu et al. proposed a data-enhanced semantic segmentation framework based on the diffusion model [104], which generates pseudo-labeled data through the diffusion model, and uses ControlNet to achieve cross-modal generation of “label→multispectral image”, and multi-stage training, which improves the segmentation accuracy.
IV-B5 Target Detection
@@
Target detection differs from pixel-level ground classification in that target detection for hyperspectral images uses bounding boxes to localize instances like an airplane, a vehicle, or a ship in the image [121, 122]. One of the most challenging problems in target detection is that the training data is ambiguous, disturbed by external conditions [123, 118], and lacks realism and diversity. This problem is due to the intrinsic nature of hyperspectral images, which usually results in objects that need to be detected being complex, difficult to separate, and distributed in different regions of the image. Therefore, augmenting the objects to be detected with diffusion models has become an effective solution.
In order to solve all the above mentioned problems, Yu et al. proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior), a diffusion model based approach [118], which introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly coordinated and lifelike mosaic images, using a powerful diffusion model prior that resamples the pixels with samples to introduce real-world information. This resampling makes the generated samples realistic and diverse. Similarly, Qi et al. proposed a novel joint framework for underwater hyperspectral image restoration and target detection based on conditional diffusion framework [123], which uses a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra, achieve high-quality restored images and superior detection performance, and effectively overcome the challenges of underwater hyperspectral imaging. spectral imaging challenges.Zhu et al. proposed ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models), a novel approach [124] to generate high-quality images conditioned on bounding boxes, thereby facilitating data synthesis for object detection.The use of synthetic visual cues with spatial constraints and textual descriptions by object to control the diffusion model allows ODGEN to perform well at demonstrate robustness when dealing with complex scenes and specific domains. Meanwhile, in order to improve the precision and accuracy, Wang et al. proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125], which combines oriented frame representation, feature alignment and accelerated diffusion, a process that detects objects directly from a set of random frames. In the training phase, the horizontal detection boxes are first converted to oriented detection boxes. The model then learns to reverse this conversion process by diffusing from ground truth oriented boxes to a random distribution. In the inference phase, the model progressively refines a set of randomly generated boxes to generate the final output.
@@ IV-B6 Anomaly Detection
Image anomaly detection is an important topic in computer vision, aiming to highlight anomalous regions of an image [115, 126, 127, 128]. In recent years, generative modeling has greatly influenced the queries for anomaly detection and shown promising results. Therefore, it will be explored in this section how the diffusion model, as the main generative model in anomaly detection, can be applied in hyperspectral collar anomaly detection. In fact, there are relatively few studies on hyperspectral image anomaly detection based on diffusion modeling, however, there are relatively numerous applications in fields such as medical impact areas [129, 130], industrial images [131, 132], etc., so with the development of diffusion modeling in this hyperspectral image anomaly detection will be increasingly used.
@@
Meanwhile, in the field of anomaly detection, it is often faced that on the one hand, the complexity of the background of the input hyperspectral image leads to the confusion of anomalies in the background. On the other hand, the lack of labeled samples for hyperspectral images leads to poor generalization of existing anomaly detection methods. To solve this problem, there Ma et al. proposed a novel solution for anomaly detection, BSDM (Background Suppression Diffusion Model) [133], which simultaneously learns the Latent Background distribution and generalizes it to different datasets to suppress the complex background. It is characterized by three aspects:(1) Pseudo background noise is designed for the complex background of HSIs and Diffusion Model (DM) is used to learn the latent background distribution therein. (2) For the generalizability problem, a statistical bias module is applied to adapt the BSDM to datasets with different domains without labeling the samples. (3) To achieve background suppression, the inference process of DM was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as noise. They were also the first to apply diffusion modeling to hyperspectral image anomaly detection. Wu et al. proposed the diffusion background dictionary method (DBD) [134], which combines the advantages of model-driven and data-driven based on the characteristics of HSI and HAD tasks.DM is intrinsically combined with the low-rank representation (LRR) model, and DM is utilized to obtain the key background dictionary tensor in the tensor LRR. Thus, anomalies can be accurately detected. According to the idea of RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to make it more suitable for suppressing the background. Also hyperspectral image anomaly detection is often faced with the lack of a priori information, essentially the challenge is to accurately reconstruct the background region while inferring the latent background of the anomalous region, due to which Chen et al. proposed a novel approach called Double Window Spectral Diffusion (DWSDiff) [56], which modeled accurate background estimation through an iterative spectral diffusion and reverse diffusion process. A dual-window strategy is also introduced to mitigate the effect of anomalous extended regions in the neighborhood on the background estimation. An anomaly generation strategy based on Principal Component Analysis (PCA) and Linear Spectral Mixture Model (LSMM) is devised to create realistic training data that overcomes the scarcity of labeled hyperspectral data. The same effect is improved.
@@
In conclusion, the road of diffusion model in hyperspectral image anomaly detection is still long and the research prospect is very great, more and more researchers will put their eyes here to promote the vigorous development of the field of anomaly detection.
IV-B7 Noise Suppression
@@
The main challenge in hyperspectral imagery is to acquire images without losing important information. During the acquisition and/or further processing stages, the acquired image may be corrupted by noise or artifacts. Noise degrades the image quality and is especially noticeable when the objects are small and the contrast is relatively low. Due to the nature of the generative model, the diffusion model facilitates the solution of various noise reduction problems [135, 136, 137]. Among them, [136] is denoising in image restoration, where the robustness of denoising is enhanced by designing a multi-shape spatial rectangular self-attention module, a spectral latent diffusion enhancement module, and also by utilizing the diffusion model, which allows for a tighter alignment with the desired HSI.
@@
In image processing, hyperspectral images are always accompanied by various noises due to the inherent limitations of imaging techniques and environmental conditions. In essence, the learning process of diffusion models is equivalent to the denoising process [19], which makes it possible to achieve excellent denoising performance in hyperspectral contexts. Different image modalities face different noise challenges. For example, hyperspectral image (HSI) recovery, which may be quite susceptible to distributional biases, is data-poor [138]. HSI needs to take into account the inhomogeneous distribution of noise in the spectral dimension [139] as well as the correlation of noise between different frequency bands [140, 141]. Fortunately, these noises can be efficiently removed by a diffusion model.Ceng and Cao et al. proposed Diff-Unmix [142] to address the challenge of recovering noise-degraded HSI through the fusion of spectral unmixing and conditional abundance generation, which employs a learnable block-based spectral unmixing strategy supplemented by a purely transformer-based backbone. We then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks. This network reconstructs noise-free decomposition probability distributions, effectively mitigating noise-induced degradation within these components. Finally, the reconstructed HSI is reconstructed by de-mixing reconstruction by mixing the diffusion-adjusted abundance maps with the spectral end-elements. he et al [140] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process instead of the pure noise in order to avoid destroying the inherently valid information in the HSI. In addition, Deng et al. proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], which improved DM-based HSI denoising by combining RGB-DM pre-training on additional RGB images, and a fusion operator was introduced to integrate HSI-DM and RGB-DM, which improves the performance of DM-based HSI denoising.Kai et al. proposed a multi-scale spatial spectral for hyperspectral image frequency joint denoising network (MSFDN) [137], in order to describe and extract the noise more accurately and obtain better denoising effect, they constructed a subspace projection network to convert the hyperspectral image into a low-dimensional subspace, designed a noise separation network, a subspace inverse projection network and utilized the conditional diffusion model for better robustness.

### v2

@@ 4.2.2 Data Generation and Completion
To address this, Zhang et al. [91] introduced a spectral folding technique that converts the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al. [92] utilized a Conditional Vector Quantization Generative Adversarial Network (VQGAN) [93] to derive a latent space for the hyperspectral map within the Latent Diffusion Model. Diffusion model training and sampling are then performed within this more manageable latent space.Furthermore, to tackle the issues of high cost and low spatial resolution, Deng et al. proposed DiffUn [94], a Diffusion Model-based hyperspectral spectral unmixing method. This approach combines spectral prior distributions learned by an unconditional Diffusion Model with likelihoods derived from the Linear Mixing Model (LMM) to enhance deblurring performance and synthesize hyperspectral maps.More recently, Dong et al. introduced a novel ISPDiff [95] model. This transparent, unified, and scale-flexible model addresses these problems using an iterative optimization algorithm. It incorporates a Multi-Scale Residual Network (MSRN) for scale-matching recovery and a Cross-Modal UNet (CMUN) for cross-scale upsampling.
@@
Overall, image-to-image generation methods based on conditional diffusion models [67, 68] fundamentally generate images from noise, rather than directly converting existing ones, by using bootstrap images as conditional inputs to conditional DDPMs. To achieve a true image-to-image conversion, Yu et al. proposed UnmixDiff [96]. This model explores the feasibility of depth generative models in HSI (Hyperspectral Image) synthesis, synthesizing HSIs from the perspective of material distribution and introducing a clear physical interpretation into the task. By integrating an unmixing autoencoder within the diffusion generation model, hyperspectral image synthesis is projected from the high-dimensional spectral domain to a low-dimensional abundance domain. This significantly reduces computational complexity and effectively boosts the spectral fidelity of the generated HSIs. Constructing a diffusion generation model in the abundance space also allows for the generation of HSIs with realistic spatial details while maintaining spectral consistency. This type of approach notably employs self-supervised learning to extract feature embeddings from input patches. These embeddings then serve as conditional inputs, guiding the diffusion model’s learning process. Crucially, this method preserves the spatial arrangement of each color patch from the original image, enabling the generated patches to be seamlessly assembled into a large and coherent image.
@@ 4.2.3 Classification
Among applications in the hyperspectral field, the classification task is paramount for image analysis, enabling accurate identification and characterization of diverse structures and anomalies within an image. It significantly aids professionals in interpreting large volumes of complex data. Despite this potential, image classification aims to extract more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98]. However, hyperspectral images inherently contain highly diverse and complex scenes and information, which complicates accurate classification. Therefore, leveraging diffusion models to enhance classification remains a significant challenge that requires further strengthening and resolution. Nevertheless, because diffusion models can learn and simulate more intricate data distributions than other deep learning models, there’s a growing commitment to applying them to hyperspectral image classification, leading to increasingly promising results [99, 100, 101, 102, 103, 104].
For image classification, the complex data distributions inherent in hyperspectral images present greater challenges for traditional deep learning models. Fortunately, diffusion models excel at capturing the joint spectral-spatial features of these images [105, 106, 107], leading to improved classification accuracy [108, 109] and enhanced generalization ability [110].
Building on these strengths, Zhou et al. [111] utilized the temporal information of diffusion models to construct a time-stepping feature library. They also introduced a dynamic fusion module to combine spectral-spatial features with temporal features, allowing for sufficient image information before classification. Li et al. [112] designed a two-branched diffusion model that extracts features from both HSI and LiDAR images, achieving information complementarity across different modalities. This approach boosts pixel discriminability and demonstrates excellent performance in more complex hyperspectral tasks [37]. Qu et al. [113] further deepened this by configuring the encoder of their two-branch diffusion model to operate in a parameter-sharing mode, ensuring shared features are extracted from multimodal images. Additionally, Chen et al. [114] leveraged diffusion models to assist in deep subspace construction, achieving impressive unsupervised HSI classification performance.
Taking a different approach, Qu et al. [115] combined a diffusion model with both a super-resolution (SR) network and a classification network. This allowed for the progressive reconstruction of high-quality images by iteratively performing image super-resolution and classification tasks, creating a mutual guidance mechanism. By utilizing multi-scale classification results to guide the target-scale image in learning new knowledge and further reconstructing high-quality images, both super-resolution and classification performance were significantly improved. Jiang et al. proposed an image fusion classification network based on Denoising Diffusion Probabilistic Models (DDPMs) [107], which demonstrated a significant improvement in classification accuracy after fusion. Similarly, Zhu et al. introduced a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN), also based on DDPMs [102]. This network incorporates a spatially self-attentive denoising module (SSAD) and a spectral group self-attentive denoising module (SGSAD). It enhances the efficiency of spectral-spatial feature learning by designing new DDPM loss functions and Contrastive Learning (CL) with Log Absolute Error (LAE). Furthermore, the inclusion of an adaptive weighted addition module (AWAM) and a cross-time step spectral spatial fusion module (CTSSFM) effectively mitigates interference from spatial spectral heterogeneity and noise effects, leading to better classification.
@@ 4.2.4 Segmentation
Image segmentation is a crucial task in computer vision, encompassing areas like semantic segmentation [116] and general image segmentation [14, 117]. In deep learning, semantic segmentation is considered a subtask of image segmentation [112, 118]. Hyperspectral image segmentation specifically aims to simplify image complexity by breaking it down into multiple meaningful segments. This process provides valuable information for hyperspectral-related fields, thereby facilitating image analysis. However, deep learning models typically demand a large volume of diverse pixel-annotated training data to yield generalizable results. Unfortunately, the number of images and labels available for hyperspectral image segmentation is limited due to the significant time, cost, and expertise required for annotation. Consequently, diffusion modeling has emerged as a promising avenue in image segmentation research, as it can synthesize labeled data, thereby eliminating the need for extensive pixel annotation [102, 104].
@@
Dong et al. [95] are utilizing Denoising Diffusion Probabilistic Models in conjunction with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks. Chen et al. [119] are designing image conditional mixing guides and integrating them with Denoising Diffusion Probabilistic Models. For labeling information, Zhang et al. [110] leverage Diffusion Probabilistic Models, while Zhu et al. [120] employ Variational Auto-Encoders alongside semi-supervised diffusion models for annotation. Li et al. proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to enhance pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism. Zhu et al. also proposed DiffCRNs [102], a diffusion-based model coupled with contrastive learning, to extract spatial-spectral features and ultimately generate pixel-level classification circles. Furthermore, Xu et al. introduced a data-enhanced semantic segmentation framework based on the diffusion model [104]. This framework generates pseudo-labeled data via the diffusion model, uses ControlNet for cross-modal "label→multispectral image" generation, and employs multi-stage training to improve segmentation accuracy.
4.2.5 Target Detection
@@
In contrast to pixel-level ground classification, target detection for hyperspectral images focuses on localizing specific instances, such as an airplane, vehicle, or ship, within the image using bounding boxes [121, 122]. A significant challenge in this area is the ambiguity and lack of realism and diversity in training data, often exacerbated by external disturbances [123, 118]. This issue stems from the inherent nature of hyperspectral images, where objects needing detection are frequently complex, difficult to distinguish, and distributed across various image regions. Consequently, augmenting the objects to be detected using diffusion models has emerged as an effective solution.
To address the aforementioned challenges, Yu et al. proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior) [118]. This diffusion model-based approach introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly realistic and diverse mosaic images. It leverages a powerful diffusion model prior that resamples pixels with samples, thereby introducing real-world information and ensuring the generated samples are both realistic and varied.
Similarly, Qi et al. presented a novel joint framework for underwater hyperspectral image restoration and target detection, based on a conditional diffusion framework [123]. This framework employs a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra. This achieves high-quality restored images and superior detection performance, effectively overcoming the inherent challenges of underwater hyperspectral imaging. Zhu et al. introduced ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models) [124], a novel approach to generate high-quality, bounding-box-conditioned images. This facilitates data synthesis specifically for object detection. By utilizing synthetic visual cues with spatial constraints and textual descriptions per object to control the diffusion model, ODGEN demonstrates robust performance when handling complex scenes and specific domains.
Meanwhile, to enhance precision and accuracy, Wang et al. proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125]. This approach combines oriented frame representation, feature alignment, and accelerated diffusion, allowing for direct object detection from a set of random frames. During the training phase, horizontal detection boxes are initially converted to oriented detection boxes. The model then learns to reverse this conversion by diffusing from ground truth oriented boxes to a random distribution. In the inference phase, the model progressively refines a set of randomly generated boxes to produce the final output.
@@ 4.2.6 Anomaly Detection
Image anomaly detection, a critical area in computer vision, focuses on highlighting unusual regions within an image [115, 126, 127, 128]. In recent years, generative modeling has significantly influenced anomaly detection research, showing promising results. This section will explore how diffusion models, as a prominent generative model, can be applied to hyperspectral image anomaly detection. While studies specifically on diffusion model-based hyperspectral image anomaly detection are relatively few, there are numerous applications in other fields, such as medical imaging [129, 130] and industrial inspection [131, 132]. Given the ongoing advancements in diffusion modeling, its use in hyperspectral image anomaly detection is expected to grow significantly in the future.
In the field of anomaly detection, a persistent challenge is the complexity of background clutter in hyperspectral images, which often obscures anomalies. Additionally, the scarcity of labeled hyperspectral samples leads to poor generalization in existing anomaly detection methods.
To tackle these issues, Ma et al. proposedBSDM (Background Suppression Diffusion Model) [133], a novel anomaly detection solution. BSDM simultaneously learns the latent background distribution and generalizes it across different datasets to suppress complex backgrounds. This approach is characterized by three key aspects: 1. Pseudo background noise is designed for the complex backgrounds of HSIs, and a Diffusion Model (DM) is employed to learn the underlying latent background distribution. 2. To address generalizability, a statistical bias module adapts BSDM to datasets from various domains without requiring labeled samples. 3. For effective background suppression, the DM’s inference process was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as if it were noise. This work also marked the first application of diffusion modeling to hyperspectral image anomaly detection.
@@
Wu et al. introduced the diffusion background dictionary method (DBD) [134], combining the strengths of model-driven and data-driven approaches based on HSI and HAD (Hyperspectral Anomaly Detection) task characteristics. DBD intrinsically integrates the DM with the low-rank representation (LRR) model, utilizing the DM to obtain the crucial background dictionary tensor within the tensor LRR, thus enabling accurate anomaly detection. Following the principles of the RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to better suit background suppression. Furthermore, hyperspectral image anomaly detection often grapples with a lack of prior information, primarily the challenge of accurately reconstructing background regions while inferring the latent background of anomalous areas. To this end, Chen et al. proposed Double Window Spectral Diffusion (DWSDiff) [56]. This novel approach models accurate background estimation through an iterative spectral diffusion and reverse diffusion process. A dual-window strategy is also introduced to mitigate the impact of extended anomalous regions in the neighborhood on background estimation. Finally, an anomaly generation strategy based on Principal Component Analysis (PCA) and the Linear Spectral Mixture Model (LSMM) is devised to create realistic training data, overcoming the scarcity of labeled hyperspectral data and leading to improved performance.
@@
In conclusion, the application of diffusion models in hyperspectral image anomaly detection remains a rapidly evolving field with significant research potential. We anticipate a growing number of researchers will focus on this area, driving substantial advancements in anomaly detection.
4.2.7 Noise Suppression
@@
A primary challenge in hyperspectral imaging is acquiring images without losing crucial information. During acquisition or subsequent processing, images can be corrupted by noise or artifacts, significantly degrading quality, especially with small objects or low contrast. Due to their generative nature, diffusion models are well-suited to addressing various noise reduction problems [135, 136, 137]. For instance, [136] focuses on denoising in image restoration by designing a multi-shape spatial rectangular self-attention module and a spectral latent diffusion enhancement module. Leveraging the diffusion model allows for a tighter alignment with the desired HSI, enhancing denoising robustness.
In image processing, hyperspectral images (HSIs) are invariably affected by various types of noise due to the inherent limitations of imaging techniques and environmental factors. Fundamentally, the learning process of diffusion models is akin to a denoising process [19], making them exceptionally effective for noise reduction in hyperspectral contexts. Different image modalities face unique noise challenges. For instance, HSI recovery can be particularly susceptible to distributional biases and often suffers from data scarcity [138]. HSIs also contend with the inhomogeneous distribution of noise along the spectral dimension [139] and the correlation of noise across different frequency bands [140, 141]. Fortunately, diffusion models can efficiently address these noise issues.
@@
Ceng and Cao et al. proposed Diff-Unmix [142] to tackle the challenge of recovering noise-degraded HSIs by fusing spectral unmixing with conditional abundance generation. This method employs a learnable block-based spectral unmixing strategy, supported by a purely transformer-based backbone. They then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks, reconstructing noise-free decomposition probability distributions and effectively mitigating noise-induced degradation within these components. Finally, the HSI is reconstructed through de-mixing by combining the diffusion-adjusted abundance maps with the spectral end-members. He et al. [140] introduced a truncated diffusion model that initiates denoising from an intermediate step of the diffusion process, rather than from pure noise. This approach aims to preserve the inherently valid information within the HSI. Additionally, Deng et al. proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], improving DM-based HSI denoising by integrating RGB-DM pre-training on supplementary RGB images. They also introduced a fusion operator to combine HSI-DM and RGB-DM, further enhancing the performance of DM-based HSI denoising. Kai et al. proposed a multi-scale spatial spectral joint denoising network for hyperspectral images (MSFDN) [137]. To more accurately describe and extract noise for superior denoising, they constructed a subspace projection network to convert HSIs into a low-dimensional subspace. They also designed a noise separation network and a subspace inverse projection network, leveraging a conditional diffusion model for enhanced robustness.